OpenStudy (anonymous):

\[\sum_{k=1}{\infty} \frac{k+5}{5^{k}}\]

OpenStudy (anonymous):

\sum_{k=1}^{\infty} \frac{k+5}{5^{k}}

OpenStudy (anonymous):

\[\sum_{k=1}^{\infty} \frac{k+5}{5^{k}}\]

OpenStudy (anonymous):

convergence or divergence problem

OpenStudy (kinggeorge):

I imagine the ratio test would work well.

OpenStudy (anonymous):

ok i'll try it

OpenStudy (anonymous):

\[ \frac{k+1+5}{5^{k+1}}*\frac{5^{k}}{k+5}\] \[ \frac{5^{k}}{5^{k+1}}*\frac{k+6}{k+5}\] \[ 5*\frac{k+6}{k+5}\]

OpenStudy (kinggeorge):

Shouldn't that be\[\frac{k+6}{5(k+5)}\]

OpenStudy (anonymous):

i don't think so, but i could be wrong

OpenStudy (anonymous):

k-k-1 is -1 so yes you are right

OpenStudy (anonymous):

so since the denominator is greater is it divergent?

OpenStudy (kinggeorge):

Careful. To use the ratio test, we need to look at \[\lim_{k\rightarrow\infty} \frac{k+6}{5k+25}=L\]If \(L<1\) it's convergent, if \(L>1\) it diverges, if \(L=1\), it tells us nothing. What is \[\lim_{k\rightarrow\infty} \frac{k+6}{5k+25}?\]

OpenStudy (anonymous):

if we multiply the numerator and dinominator by 1/k and let k-> infty we get 6/25

OpenStudy (kinggeorge):

I believe it would be 1/5 and not 6/25, but in either case, this is less than 1, so the series converges.

OpenStudy (anonymous):

i think you're right....but why 1/5 sir George

OpenStudy (kinggeorge):

\[\lim_{k\rightarrow\infty} \frac{k+6}{5k+25}\cdot \frac{1/k}{1/k}=\lim_{k\rightarrow\infty} \frac{1+6/k}{5+25/k}\]Since \(6/k\) and \(25/k\) go to 0, we get 1/5.

OpenStudy (anonymous):

i'm soo silly thanks

OpenStudy (kinggeorge):

you're welcome.

OpenStudy (anonymous):

it's late where i'm from...i'm forgetting basic algebra

OpenStudy (kinggeorge):

I've done much worse. For example, I could have sworn 9+4=11 not too long ago.

OpenStudy (anonymous):

lolz it happens to the best of us :)